Christie purchases a one-third interest in the Corporate Capital Partnership (CCP) in 2018 for $44,900. During 2018, CCP earns an income of $80,700, and Christie withdraws $31,300 in cash from the partnership. In 2019, CCP suffers a loss of $24,000, and Christie withdraws $9,600. What are the tax consequences for Christie of this investment in 2018 and 2019?

Christie (will or will not include) her share of the partnership income, $_____ in 2018 and deduct her share of the partnership _______ , $ in 2019.
Answer #1

Christie will include her share of the partnership income in 2018 = 80700 * 1/3 = 26900

Christie will deduct her share of the partnership in 2019 = 24,000 * 1/3 = 8000